# Collaborative Chain-of-Agents for Parametric-Retrieved Knowledge Synergy

### v2

@@ 3.5 Ablation Study I: Different Modules
Table 3: Ablation study on knowledge induction and decision-making. The zero-shot variant (§ 2.3) is also included. We use the average of EM and F1 for fair evaluation.
3.6 Ablation Study II: Training Strategies
@@
Table 4: Ablation study of the training strategy for CoCoA. For fairness, the average of EM and F1 is used as the metric.
@@ 3.7 When the Number of K Changes
3.8 Performance of Different Model Sizes
Figure 5: Illustration of accuracy changes at different model sizes, with Avg(EM,F1) as the metric.
3.9 Training Generalization to Non-QA Tasks
To further evaluate the generalization ability of CoCoA, we test its performance on fact verification and multiple-choice tasks. As shown in Figure  6, our training did not reduce the performance of these tasks compared to standard RAG. In fact, in some cases, we even observed a slight improvement. One explanation is that our training strategy encourages collaborative output that leverages the capabilities of the LLM, rather than injecting knowledge directly, and thus possesses a certain degree of universality.
4 Related Works
4.1 Retrieval-augmented Generation
In recent years, to address outdated knowledge and hallucination of LLM, RAG has been introduced (Fan et al. 2024; Gao et al. 2023), and many efforts have been made in two aspects: “how to retrieve more relevant information” including retriever fine-tuning (Nian et al. 2024) and query optimization (Ma et al. 2023; Wang, Yang, and Wei 2023; Wang et al. 2024b) and “how to better use the retrieved information” including domain fine-tuning (Wang et al. 2024c; Zhang et al. 2024; Yue et al. 2025; Xia et al. 2025) and controlled decoding strategies (Shi et al. 2023). Our CoCoA falls into the second category: better utilization of knowledge.
4.2 RAG Pipeline Optimization
Pipeline optimization usually adds pre-generation processing, retrieval intent identification, or optimizes the pipeline as a whole. For example, Glass et al. (2022); Kim and Lee (2024) and Yu et al. (2023) introduce reranking and refinement steps before generation, mitigating the impact of noisy retrieved passages. SKR (Wang et al. 2023) and UAR (Cheng et al. 2024) avoid unnecessary retrieval by adding retrieval intent identification processes before generation. SURE(Kim et al. 2024a) first generates multiple candidate answers and performs conditional summary verification based on the candidate answers, allowing LLMs to focus on specific contexts. However, these methods either rely too much on retrieved content or fail to combine internal and external knowledge, which can limit performance.
4.3 RALM Enhancement
Retrieved-Augmented Language Model(RALM) enhancement is usually achieved by adjusting the LLM to achieve effective use of the information. One common approach is to train the LLM itself. For example, RAFT (Zhang et al. 2024) and InstructRAG (Wei, Chen, and Meng 2024) improve the model’s ability to resist noise in external context by introducing noise resistance training. REAR (Wang et al. 2024c) achieves the model’s trade-off between external context and internal knowledge by training the model’s relevance-guided generation capabilities. Self-RAG (Asai et al. 2023) trains LLMs to decide whether to perform retrieval and to improve their self-reflection ability. Another approach involves guiding the decoding (Shi et al. 2023; Kim et al. 2024b). For instance, CAD (Shi et al. 2023) enforces absolute trust in retrieved information by using contrastive decoding under the assumption that external information is fully correct. However, both approaches tend to underutilize the model’s internal knowledge, which may constrain the quality and informativeness of its responses.
@@ 5 Conclusion
We investigate the challenge of parametric–retrieved knowledge collaboration and introduce CoCoA, a retrieval-augmented generation framework that improves LLM performance. By leveraging a two-stage multi-agent pipeline, CoCoA-zero collaborates internal and external knowledge and provides self-synthesized supervisory signals. Based on a long-chain training strategy, CoCoA delivers strong results on QA tasks, demonstrating its effectiveness and offering insights into long-chain reasoning and collaborative agent training for knowledge-intensive applications.
@@ Appendix A Dataset
2WikiMultiHopQA (Ho et al. 2020b) and HotpotQA (Ho et al. 2020a): Both datasets are multi-hop question answering datasets based on Wikipedia. Considering the limitation of experimental cost, we used the sub-sampling set published by Trivedi et al. (2022); Kim et al. (2024a), which is obtained by extracting 500 questions from the validation set of each dataset.
WebQuestions (Berant et al. 2013): Constructed from questions posed by the Google Suggest API, where the answers are specific entities listed in Freebase.
TriviaQA (Joshi et al. 2017): A compilation of trivia questions paired with answers, both originally pulled from online sources.
@@
We sampled subsets from the training sets of HotpotQA (Ho et al. 2020a), 2WikiMultiHopQA (Ho et al. 2020b) and WebQuestions (Berant et al. 2013), then used the CoCoA-zero framework to synthesize data and filtered them with gold answers. Finally, we selected 6.8k filtered samples, including 3k, 3k, and 0.8k from the three datasets, respectively. For the DPO training data, we screen out 1151 samples, which are the ones that are answered incorrectly by zero-shot but correctly by the CoCoA-zero. For each sample, we gathered 5 relevant passages using the most common retriever Contriever (Izacard et al. 2021).
Appendix B Baseline Setting
We followed the original settings for almost all experiments. For baselines requiring training, we directly used their weights. Note that InstructRAG directly generates long rationales, the first half of which consists mostly of analysis and citations of the document, resulting in a non-strictly high EM score and a low F1 score. For a fair comparison, we used Qwen2.5-3B to perform answer segmentation to evaluate.
@@ Appendix C Training Details
We fine-tune LLaMA3.1-8B with LoRA (r=16, $\alpha$=16, dropout=0.05) on a maximum input length of 2048. LoRA is applied to attention projection layers. During SFT, we trained for 5 epochs with a batch size of 1, gradient accumulation of 4, and a learning rate of 3e-5. For DPO, a $\beta$ value of 0.2 is applied, using a sigmoid loss function, while RPO is configured with an $\alpha$ value of 0.2. The learning rate was set to 5e-6 and other settings are the same as SFT. During inference, we use the vllm (Kwon et al. 2023) accelerated inference framework, and to ensure repeatability, we set the temperature to 0.0. All experiments are conducted on a single A100 GPU with 80GB or 40GB memory.
@@
Appendix D Optimization Analysis
@@
$\frac{\partial}{\partial\theta}\mathcal{L}_{\text{indep}}$ $=\frac{\partial}{\partial\theta}\ \bigl{[}-\log P_{\theta}(s\mid x,d)\bigr{]}$ (14)
@@
$\frac{\partial}{\partial\theta}\mathcal{L}_{\text{chain}}$ $=\frac{\partial}{\partial\theta}\bigl{[}-\log P_{\theta}(s\mid x,d)\bigr{]}+\frac{\partial}{\partial\theta}\bigl{[}-\log P_{\theta}(\hat{a}\mid x,s,d)\bigr{]}$ (15)
@@
Here, $\Delta_{g}:=\frac{\partial}{\partial\theta}\bigl{[}-\log P_{\theta}(\hat{a}|x,s,d)\bigr{]}$ is the additional gradient that the answer‑loss naturally back‑propagates to the pre-processing parameters when the same network $\theta$ produces both tokens. In the independent setting $\Delta_{g}=0$ by construction, so the preprocessor never “hears” whether the answer is correct, which is not conducive to the consistency of the response. The chain objective restores this missing credit assignment signal, thus performing a special kind of multi-task learning on both stages, optimizing them instead of each in isolation, potentially helping to escape from local optimal solutions.
Appendix E Full Results
@@
Appendix F Prompt Templates
@@
PubHealth Is the following statement correct or not? Say true if it’s correct; otherwise say false. Don’t capitalize or add periods, just say ”true” or ”false”. Do not exceed one word.
@@
Appendix G Limitations

### v3

@@ 4.5 Ablation Study I: Different Agents
Table 2: Ablation study on knowledge induction and decision-making. The zero-shot variant (§ 3.3.2) is also included. We adopt the EM/F1 average as metric.
4.6 What is the Impact of Internal Conditional Induction?
To explore whether internal knowledge generation can introduce harm, we conducted a qualitative analysis of our sample. Specifically, we analyzed (1) answers generated using internal knowledge and (2) answers generated directly, as shown in Figure 3. We found that internal knowledge generation had some failures, but successes far outweighed failures. This suggests that conditional internal knowledge induction is a knowledge induction method and has a certain secondary validation effect, but it inevitably introduces some adverse effects such as hallucinations, which require more fine-grained control in the future.
Figure 3: Illustration of Conditional Induction, with EM as the metric.
Figure 4: Qualitative Case Analysis (EM Metric): EW-IW (both win), EL-IL (both lose), EW-IL (external wins), EL-IW (internal wins). The outer loop is the direct answer of the internal and external agents, and the inner loop is CoCoA-zero.
4.7 Qualitative Case Study
To analyze the mechanism of CoCoA-zero’s effectiveness, we conducted a detailed qualitative analysis of the cases. The outer loop represents the proportion of responses directly generated by internal and external responses. We find that both achieved success and failed responses, demonstrating a contest of knowledge. The inner loop represents CoCoA-zero responses. We find that CoCoA-zero effectively handles cases where both responses are correct, achieving coverage rates of 57%+ on 2WikiMQA and 73%+ on HotpotQA, surpassing the coverage of individual responses. This is inconsistent with the conclusion that some work LLM prefers internally generated content(Tan et al., 2024). We speculate that conditional induction may alleviate this bias.
Furthermore, we find that even when both responses failed, CoCoA-zero also achieved some success. This stems from the mutual enhancement between knowledge induction and high-level decision-making, which jointly enable the LLM to better activate its own capabilities.
@@ 4.7 Qualitative Case Study
Table 3: Ablation study of the training strategy for CoCoA. For fairness, Avg(EM,F1) is used as the metric.
4.8 Ablation Study II: Training Strategies
4.9 Training Generalization to Non-QA Tasks
To further evaluate the generalization of CoCoA, we test it on fact verification and multiple-choice tasks. As shown in Figure  5, our training did not reduce the performance of these tasks compared to standard RAG. In fact, in some cases, we even observed a slight improvement. One explanation is that our training strategy encourages collaborative output that leverages the capabilities of the LLM, rather than injecting knowledge directly, and thus possesses a certain degree of universality.
@@ 4.10 When the Number of K Changes
Figure 6: Performance varies with passage num: 2WikiMQA (left), WebQuestions (right).
4.11 Inference Efficiency
To evaluate the prospect of CoCoA in practical applications, we study its inference efficiency. As shown in Table 4, our CoCoA is three times more expensive than common methods (eg., CON), requiring a budget of about 600 output tokens. However, compared to reasoning models (eg., R1), the cost still has a significant advantage, especially when the performance is improved. This achieves a trade-off between performance and efficiency.
Method Avg. Input Avg. Output Avg(EM,F1)
CoT 904.184 195.882 23.83
R1-8B 829.186 818.200 31.30
CoCoA 993.184 609.742 41.29
Table 4: Efficiency analysis on 2WikiMQA, with average token count reported.
@@ 5 Conclusion
We investigate the challenge of parametric–retrieved knowledge synergy and introduce CoCoA, a RAG framework that improves LLM performance. By leveraging a two-stage multi-agent pipeline, CoCoA-zero integrates internal and external knowledge and provides self-synthesized supervisory signals. With long-chain training, CoCoA delivers strong results on QA tasks, demonstrating its effectiveness and offering insights into long-chain reasoning and collaborative agent training for knowledge-intensive applications.
Limitations
@@ Appendix A Dataset
2WikiMultiHopQA (Ho et al., 2020b) and HotpotQA (Ho et al., 2020a): Both datasets are multi-hop question answering datasets based on Wikipedia. Considering the limitation of experimental cost, we used the sub-sampling set published by Trivedi et al. (2022); Kim et al. (2024a), which is obtained by extracting 500 questions from the validation set of each dataset.
WebQuestions (Berant et al., 2013): Constructed from questions posed by the Google Suggest API, where the answers are specific entities listed in Freebase.
TriviaQA (Joshi et al., 2017): A compilation of trivia questions paired with answers, both originally pulled from online sources.
@@
We sampled subsets from the training sets of HotpotQA (Ho et al., 2020a), 2WikiMultiHopQA (Ho et al., 2020b) and WebQuestions (Berant et al., 2013), then used the CoCoA-zero framework to synthesize data and filtered them with gold answers. Finally, we selected 6.8k filtered samples, including 3k, 3k, and 0.8k from the three datasets, respectively. For the DPO training data, we screen out 1151 samples, which are the ones that are answered incorrectly by zero-shot but correctly by the CoCoA-zero. For each sample, we gathered 5 relevant passages using the most common retriever Contriever (Izacard et al., 2021).
Task Type Datasets # Samples
Multi-HopQA 2WikiMultiHopQA 500
HotpotQA 500
OpenQA WebQuestions 2032
TriviaQA 11313
Table 5: Description of tasks and evaluation datasets.
Appendix B Baselines
We selected several of the most representative methods for comparison.
  • StandardRAG, which is the most classic “retrieve-then-read” paradigm.
  • Chain-Of-Thought (Wei et al., 2022): Uses CoT prompting to generate reasoning steps before producing the final answer.
  • Chain-Of-Note (Yu et al., 2023): Refines the retrieved passages prior to answering.
  • GenRead (Yu et al., 2022): Generates self-contained intermediate context to answer, effectively replacing retrieval with generation.
  • SURE (Kim et al., 2024a): Conditional summarization followed by multiple validation.
  • Self-RAG (Asai et al., 2023): Employs adaptive retrieval and self-reflection to decide when and how to use external knowledge.
  • DeepSeek-R1-Distill-8B (Guo et al., 2025): A distilled LLaMA-8B model released by DeepSeek-R1, trained on reasoning data.
  • InstructRAG (Wei et al., 2024): Denoising training using self-synthesized data.
Baseline Setting.
We followed the original settings for almost all experiments. All retrieval-based methods use top-5 passages. For baselines requiring training, we directly used their weights. Note that InstructRAG directly generates long rationales, the first half of which consists mostly of analysis and citations of the document, resulting in a non-strictly high EM score and a low F1 score. For a fair comparison, we used Qwen2.5-3B to perform answer segmentation to evaluate.
@@ Appendix C Training Details
We fine-tune LLaMA3.1-8B with LoRA (r=16, $\alpha$=16, dropout=0.05) on a maximum input length of 2048. LoRA is applied to attention projection layers. During SFT, we trained for 5 epochs with a batch size of 1, gradient accumulation of 4, and a learning rate of 3e-5. For DPO, a $\beta$ value of 0.2 is applied, using a sigmoid loss function, while RPO is configured with an $\alpha$ value of 0.2. The learning rate was set to 5e-6 and other settings are the same as SFT. During inference, we use the vllm (Kwon et al., 2023) accelerated inference framework, and to ensure repeatability, we set the temperature to 0.0. All experiments are conducted on a single A100 GPU with 80GB or 40GB memory.
C.1 Performance of Different Model Sizes
Figure 7: Illustration of performance changes at different model sizes, with Avg(EM,F1) as the metric.
Appendix D Performance of Different Retriever
Figure 8: Illustration of performance changes at different retrievers, with Avg(EM,F1) as the metric.
In order to verify the robustness to different retrievers, we selected BM25, DPR, CONTRIEVER and E5 as retrievers. The experimental results are shown in Figure 8. We found that different datasets have different preferences for retrievers, each with its own advantages and disadvantages. Overall, both our CoCoA-zero and CoCoA achieve robust performance on different retrievers.
@@ Appendix D Performance of Different Retriever
Appendix E Optimization Analysis
@@
$\frac{\partial\mathcal{L}_{\text{indep}}}{\partial\theta}$ $=\frac{\partial\,\bigl[-\log P_{\theta}(s\mid x,d)\bigr]}{\partial\theta}$ (11)
@@
$\frac{\partial\mathcal{L}_{\text{chain}}}{\partial\theta}$ $=\frac{\partial\,\bigl[-\log P_{\theta}(s\mid x,d)\bigr]}{\partial\theta}$ (12)
$\qquad+\frac{\partial\,\bigl[-\log P_{\theta}(\hat{a}\mid x,s,d)\bigr]}{\partial\theta}$
@@
$\Delta_{g}\;:=\;\frac{\partial\,\bigl[-\log P_{\theta}(\hat{a}\mid x,s,d)\bigr]}{\partial\theta}.$ (13)
Here, $\Delta_{g}$ is the additional gradient that the answer‑loss naturally back‑propagates to the pre-processing parameters when the same network $\theta$ produces both tokens. In the independent setting $\Delta_{g}=0$ by construction, so the preprocessor never “hears” whether the answer is correct, which is not conducive to the consistency of the response. The chain objective restores this missing credit assignment signal, thus performing a special kind of multi-task learning on both stages, optimizing them instead of each in isolation, potentially helping to escape from local optimal solutions.
Appendix F Full Results
@@
Appendix G Prompt Templates
@@
PubHealth Is the following statement correct or not? Say true if it’s correct; otherwise say false. Don’t capitalize or add periods, just say "true" or "false". Do not exceed one word.
@@
Appendix H Case Study
We provide detailed case study in Figure 9, Figure 10, and Figure 11.
Figure 9: Direct generation is correct, while standard RAG generation is incorrect. This is due to external noise causing RAG to misjudge, while CoCoA performs well in synergy.
Figure 10: Direct generation is incorrect, while standard RAG is correct. This is due to internal hallucinations causing misjudgment, while CoCoA performs well.
Figure 11: Direct generation is incorrect, and standard RAG generation is also incorrect. CoCoA synergizes well, benefiting from the interaction between knowledge induction and high-level decision-making to stimulate model capabilities.